Skip to main contentSkip to solution

In binomial distribution with n=10n = 10 and P=13P = \frac{1}{3}, the probability of the event that unequal number of failures and successes occur is :

Solution

✅ Correct Option: 1

Successes ≠\neq failures means X≠5X \neq 5. P(X=5)=(105)(13)5(23)5=252×3259049=806459049=8966561P(X = 5) = \binom{10}{5}\left(\frac{1}{3}\right)^5\left(\frac{2}{3}\right)^5 = \frac{252 \times 32}{59049} = \frac{8064}{59049} = \frac{896}{6561}. So P(X≠5)=1−8966561=56656561P(X \neq 5) = 1 - \frac{896}{6561} = \frac{5665}{6561}.

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question